If 2 numbers have a product of 75, a quotient of 3, and a difference of 10, what's the sum of these 2 numbers?
Answer by venugopalramana(3286)

let the 2 numbers be x and y .we have
xy=75.......I..........
x/y=3.................II.
x-y=10....................III
.actually 2 relations should do to solve the problem
from II,we have....x/y=3
x=3y.......
substituting in I we get
xy=3y*y=75
y*y=75/3=25
y=5
so x =3y=3*5=15
so 5 and 15 are the 2 numbers .incidentally this satoisfies the third relation of x-y =15-5=10...so no problem this is the answer...otherwise we would have to say NO ANSWER!!!!